OpenStudy (lgbasallote):

X students took an exam on Math and Physics. The probability of passing the Mathematics exam is 70% while the probability of passing the Physics exam is 50%. If no student failed both subjects, and 8 students passed both subjects, how many students took the exam?

OpenStudy (anonymous):

between 8 and infinity

OpenStudy (lgbasallote):

......

OpenStudy (lgbasallote):

you do realize it's impossible to have infinite students right?

OpenStudy (anonymous):

on our mere mortal world. I dont think I understand.

OpenStudy (anonymous):

if there is 70% chance of passing the physics exam? then there is some small chance that say 100,000 people pass the exam? I dont understand how you can work out how many people passed.

OpenStudy (lgbasallote):

i believe a contingency table is used

OpenStudy (anonymous):

Sorry, ive not heard of that, i will google

OpenStudy (lgbasallote):

wonderful

OpenStudy (anonymous):

Is this discreet math, or a probability course?

OpenStudy (lgbasallote):

...well it's in the probability section....so i would say probability

OpenStudy (lgbasallote):

i just noticed my question doesn't have a header "Probability"

OpenStudy (anonymous):

its answer is 8

OpenStudy (anonymous):

I mean, if they're independent events, then it's just \((.80)(.70)n = 8\) and you solve for \(n\).

OpenStudy (lgbasallote):

it's not that simple..

OpenStudy (lgbasallote):

my questions never are....

OpenStudy (anonymous):

so the 70% and 50% is based on the students that took the exam this time were talking aobut

OpenStudy (anonymous):

its not a constant 70% guaranteed

OpenStudy (lgbasallote):

it's constant

OpenStudy (anonymous):

so every time this exam happens, 70% of people pass. its always 70% not just 70% this time

OpenStudy (anonymous):

You're right, it's unlikely we can assume they're independent. In fact since their complements seem mutually exclusive... they're most likely dependant. Let \(A\) be the event you pass Math, \(B\) be the event you pass physics. So no one failed both exams, meaning \[P(A^C\cap B^C)=0 \]We know what \[P(A^C\cup B^C) = P(A^C)+P(B^C)+P(A^C\cap B^C) = 0.30+0.50+0=0.80 \]

OpenStudy (lgbasallote):

...drawing a contingency table is so much easier...

OpenStudy (anonymous):

Then why didn't you do it?

OpenStudy (anonymous):

he is encouraging debate

OpenStudy (anonymous):

cool gys

OpenStudy (lgbasallote):

because i don't know how

OpenStudy (anonymous):

oh lol1

OpenStudy (lgbasallote):

that's why i want to see how to do it

OpenStudy (anonymous):

i have too much faith in you

OpenStudy (lgbasallote):

there will be time for making OpenStudy users feud.

OpenStudy (anonymous):

i still dont udnerstand why there cant have been infinate that passed one and failed the other, if the 70% is constant and doesnt just relate to this test

OpenStudy (anonymous):

\[P(A^C\cup B^C) = P((A\cap B)^C) = 1-P(A\cap B)\]

OpenStudy (lgbasallote):

becuse infinity is never accepted in probability and statistics

OpenStudy (anonymous):

anybody please solve my qyestion its very intresting

OpenStudy (lgbasallote):

hmm....it seems @wio is on to something....

OpenStudy (anonymous):

okay, well what about 99999999 students, surely any amount above 8 is possible?

OpenStudy (anonymous):

\[0.80=1-P(A\cap B) \rightarrow P(A \cap B) = 0.20\]

OpenStudy (anonymous):

Thus \[0.2 n = 8\]

OpenStudy (lgbasallote):

ahh yes. i remember now

OpenStudy (anonymous):

40 people took the test.

OpenStudy (anonymous):

40 people took the test, there was (7/10) x (5/10) = 35/100 chance of passing. so 40 x (35/100) should = 8, but it = 14?

OpenStudy (anonymous):

@JamesWolf That only works if you assume that passing the tests are independent events. That isn't the case though. If you're good at math, you'll be decent at physics and vise versa.

OpenStudy (anonymous):

right, so dont you need to know by how much being good at maths helps physics?

OpenStudy (lgbasallote):

still thinking how to set up a contingency table though

OpenStudy (anonymous):

like, what if its "slightly" independant, what if its completely not independat i.e. the same

OpenStudy (anonymous):

You mean "slightly" dependent. Things are either independent or they are not. Anyway, look at algebra.

OpenStudy (anonymous):

So those people who passed the math exam 7/10 will do better on the physics. but why dont you need to know HOW much better they will do?

OpenStudy (anonymous):

Okay, want me to explain what I did...?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

\[P(A^C\cup B^C) = P(A^C)+P(B^C)-P(A^C\cap B^C) = 0.30+0.50-0=0.80\] So remember that \(A\) is the event that you pass math, \(B\) is the event you pass physics. \(A^C\) is the event you don't pass physics. \(P(A)\) would be the probability you pass math. \(P(A^C \cap B^C)\) is the probability you don't fail math and don't fail physics. Since they say no one failed both, we assume this to be zero. \(P(A^C \cup B^C)\) is the probability you fail math or fail physics. We calculate this to be 80%. Does this make sense so far?

OpenStudy (anonymous):

hold on one sec, reading

OpenStudy (anonymous):

P(AC∩BC) is the probability you don't fail math(YOU PASS) and don't fail physics (YOU PASS. Since they say no one failed both, we assume this to be zero. but 8 passed?

OpenStudy (anonymous):

P(AC∩BC) = 0?

OpenStudy (anonymous):

hahr ight i didnt see the n and u

OpenStudy (anonymous):

sorry im not familiar with this notation

OpenStudy (anonymous):

It's set notation. \(\cap\) means 'and' while \(\cup\) means 'or' when it comes to set representing events.

OpenStudy (anonymous):

so this \[P(A^C \cap B^C)\] is probability you fail A AND fail B

OpenStudy (anonymous):

\[-P(A^C \cap B^C)\] is the probability that you pass A AND pass B ?

OpenStudy (anonymous):

how should i interpret the - sign?

OpenStudy (anonymous):

remember that probability ranges from 0 to 1 (0% to 100%). Also \(1-P(X) = P(X^C)\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Basically, the probability that something does happen OR doesn't happen is 100%

OpenStudy (anonymous):

Which is why \[P(X)+P(X^C)=1 \rightarrow P(X^C)=1-P(X)\]

OpenStudy (anonymous):

So \(-P(X)\) has no meaning... but \(1-P(X)\) does.

OpenStudy (anonymous):

okay, im not sure about this bit \[P(A^C\cup B^C) = P(A^C)+P(B^C)-P(A^C\cap B^C)\] reading it... the probability of not A or not b equals the probability of A, plus the probability of B minus the probability of A and B

OpenStudy (anonymous):

am i reading that right?

OpenStudy (anonymous):

the probability of not A or not b equals the probability of not A, plus the probability of not B minus the probability of not A and not B

OpenStudy (anonymous):

sorry that ^

OpenStudy (anonymous):

Yes, that is correct. The reason we subtract not A and not B is to get rid of overlap.

OpenStudy (anonymous):

so since the equation balances could you divide through by "not" to get the probability of A or B equals the probability of A plus the probability of B minus the probability of A and B

OpenStudy (anonymous):

what do you mean by overlap?

OpenStudy (anonymous):

Umm, yeah we could have done that. We didn't know the probability of A and B though.

OpenStudy (anonymous):

Overlap as in... suppose you have a group of twenty people. Have are guys, half are adults, and a quarter are both adult an guys. If you just add probability of guys and probability of adults together, you will not get the probability of being an adult or a guy, because you will count the people who are both guy and adult twice.

OpenStudy (anonymous):

To only count them once, you have to subtract the overlap.